if i have 67% in math and the final exam is worth 30%, what grade do i need in the final exam to get a 50 in the whole subject?
Answer by drk(1908)

70% is done and you currently have 67%.
multiply these to get 46.9
You want a 50 which means you need 3.1 more points, so we set up a proportion as
30/50 = 3.1/x
cross multiply to get
30x = 155
divide to get
x = 5.166
but
5.16666/50 = 10.33%
so, you need 10.33% on the final to get 3.1 points which would = 50.